OpenStudy (anonymous):

14x^2-686/4x^2-196 can someone help please

OpenStudy (anonymous):

\[\frac{14x^2-686}{4x^2-196}=\frac{7-686}{2-196}=\frac{-679}{-194}\]\[\frac{7}{2}\]

OpenStudy (anonymous):

thank you nancy, I think I understand this better now.

OpenStudy (anonymous):

Another way to do it is take the GCF, giving you \[14(x^2-49)\div(4(x^2-49)\] Since the top and bottom both have a (x^2-49), they can cancel. This will leave you with 14/4, which simplifies to 7/2 or 3.5

OpenStudy (anonymous):

thank you entezflare